OpenStudy (anonymous):

PLEASE HELP!!!!! Describe how to transform the graph of f into the graph of g. f(x) = x6 and g(x) = - x6

OpenStudy (anonymous):

is x6 = x^6?

OpenStudy (anonymous):

i dont know... i dont understand and if i dont pass and finish my class today i dont graduate.....

OpenStudy (anonymous):

I'm guessing you mean x^6. x^6 = x^3. -1(x^6) = -x^6. Just multiply it by -1